OpenStudy (anonymous):

Susan wants to buy a paddle boat for $840. She will pay 20% down and pay the rest in six monthly installments. What will be the amount of each monthly payment?

OpenStudy (anonymous):

Hi there! Would you like to learn how to solve this problem, or just get an answer?

OpenStudy (anonymous):

Know the answe rwould be enough thank u

OpenStudy (anonymous):

112

OpenStudy (anonymous):

I'd love a medal.

OpenStudy (anonymous):

Take 840 · .20 then 840 - 168 = 672 672 / 6 = 112 $112 a month is your answer.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

840 * 0.8 / 6 = 112